# The Impact of Astrophysical Priors on Parameter Inference for GW230529

### v1

@@ Abstract
We investigate the effects of prior selection on the inferred mass and spin parameters of the neutron star-black hole merger GW230529_181500. Specifically, we explore models motivated by astrophysical considerations, including massive binary and pulsar evolution. We examine mass and spin distributions of neutron stars constrained by radio pulsar observations, alongside black hole spin observations from previous gravitational wave detections. We show that the inferred mass distribution is highly dependent upon the spin prior. Specifically, under the most restrictive, binary stellar evolution models, we obtain narrower distributions of masses with a black hole mass of $4.1^{+0.2}_{-0.3}\,M_{\odot}$ and neutron star mass of $1.3^{+0.1}_{-0.1}\,M_{\odot}$ where, somewhat surprisingly, it is the prior on component spins which has the greatest impact on the inferred mass distributions. Re-weighting using neutron star mass and spin priors from observations of radio pulsars, with black hole spins from observations of gravitational waves, yields the black hole and the neutron star masses to be $3.8^{+0.5}_{-0.6}\,M_{\odot}$ and $1.4^{+0.2}_{-0.1}\,M_{\odot}$ respectively. The sequence of compact object formation — whether the neutron star or the black hole formed first — cannot be determined at the observed signal-to-noise ratio. However, there is no evidence that the black hole was tidally spun up.
@@ 1 Introduction
The gravitational wave (GW) event GW230529_181500 (hereafter abbreviated GW230529) was detected by the LIGO Livingston Observatory during the first part of the fourth observing run of the LIGO-Virgo-KAGRA collaboration (Abac et al., 2024). The binary components are a neutron star and another compact object whose mass likely falls within the “lower mass gap” between the heaviest neutron stars and the lightest black holes, which was previously thought to exist between $2.5\,M_{\odot}$ to approximately $5\,M_{\odot}$ (Bailyn et al., 1998; Özel et al., 2010; Farr et al., 2011). Moreover, the event marks the third confident detection of a black hole-neutron star binary through gravitational waves (Abbott et al., 2021). It followed a probable observation of a pulsar-(mass gap) black hole in the globular cluster NGC 1851 (Barr et al., 2024). The theory has speculated on the existence of black hole-neutron star binaries (O’Shaughnessy et al., 2008; Fryer et al., 2012; Chattopadhyay et al., 2021; Broekgaarden et al., 2021). Still, successive detections of such objects spark more debates on their mass-spin distributions to formation channels (Chandra et al., 2024) to possible electromagnetic counterparts (Zhu et al., 2024; Ronchini et al., 2024).
@@ 2 Incorporating Astrophysical Prior Beliefs
where the spin, $\boldsymbol{\chi}_{i}$, is defined as $\boldsymbol{\chi}_{i}=\frac{\mathbf{S}_{i}}{m_{i}^{2}}$, and the normalization is chosen such that a maximally spinning BH has $|\boldsymbol{\chi}|=1$. Consequently, the inferred values of the component masses and spins from a GW observation are correlated. Furthermore, for a low signal to noise ratio (SNR) signal such as GW230529, the choice of priors can impact the inferred properties. Indeed, due to correlations between the inferred mass ratio and effective spin, choices of spin priors can impact the inferred masses and vice versa.
The gravitational-wave estimation of these parameters for GW230529, as presented in Abac et al. (2024), was performed with minimally informed priors on the mass and spin distributions of the components of the binary, in addition to the location and orientation parameters (Veitch et al., 2015). Specifically, the analysis used mass priors that are flat in the redshifted component masses, within a prescribed range of masses and mass ratios, and uniform in spin magnitude and orientation (see Appendix D of Abac et al. (2024) for details). The output of the original analysis is a set of samples from the posterior probability distribution of the parameters (Collaboration et al., 2024) [^2] Specifically, we make use of the Combined_PHM_highSpin results for the majority our our analyses. When re-weighting using the astrophysical models described in Section 3.1, which restrict component spins to be close to zero, we use IMRPhenomPv2_NRTidalv2_lowSpin to ensure that we have sufficient samples following re-weighting . We then apply weights to these samples in the ratio of our desired astrophysical prior to the original, uniform prior. Given an astrophysically motivated prior $\pi_{A}$, the posterior distribution for the parameters $\boldsymbol{\theta}$ describing the progenitor of GW230529 is given by
@@
When performing prior re-weighting, however, it is important that the proposed prior provides support in the same part of the parameter space as the target prior, otherwise the re-weighting will be highly inefficient. Since the original priors are near-uniform in the parameters of interest and also cover a broad parameter range, the astrophysical priors generally comprise a subset of the original ranges. However, in cases where the astrophysical priors are sharply peaked, the re-weighting procedure summarized above can lead to a low number of samples in the astrophysical posterior. In the analysis presented in Section 4, we use those samples to create one-dimensional Kernel Density Estimators of the inferred masses and spins. In general one needs to compute the number of effective samples (Kish, 1995; Víctor Elvira, 2018). As 90% credible intervals are the relevant numbers for the scientific conclusions of this work, we checked via bootstrapping tests to what precision the one-dimensional credible intervals from Section 4 could be quoted, and used that precision.
@@ 3.2 Models motivated by radio pulsar observations in star clusters
As an alternative approach to modelling isolated binary evolution, here we utilize the observed population of Galactic globular cluster radio millisecond pulsar [^3] Over a third of all observed millisecond pulsars appear in Galactic globular clusters, which collectively account for less than 0.05% of the total number of stars in the Milky Way spin distribution from the Australia Telescope National Facility (ATNF) pulsar catalogue [^4] https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester et al., 2005). From the catalogue, we obtain the pulsar’s spin $P$, and get
@@
where, $c$ is the velocity of light, $G$ is the universal constant of gravitation, $M$ is the mass of the pulsar (assumed to be 1.4 $M_{\odot}$ here) and $I$ is the moment of inertia of the pulsar (computed from Lattimer & Schutz, 2005, with mass=1.4 $M_{\odot}$ and radius$=12$ Km). Although most pulsars are expected to spin down significantly due to the loss of rotational energy, with decaying magnetic field at the time of merger (Figure 17, 21 of Chattopadhyay et al., 2021), we take the observed NS spin distribution as an upper limit. The maximum observed spin $\chi_{\mathrm{NS}}$ is 0.4, with the primary peak at 0.13 and another at about 0.02. The mass distribution for NSs, while dependent on the NS equation of state and the assumed Tolman-Oppenheimer-Volkoff limit for theoretical studies, can also be estimated from pulsar observations (with the inclusion of radio selection effects). Even with the limited dataset due to the difficulty in measuring NS mass, there have been multiple studies to determine the shape and range of the mass distribution of NSs (Antoniadis et al., 2016; Alsing et al., 2018; Rocha et al., 2023). For this analysis, we assume the NS mass distribution to be a double Gaussian with a maximum mass of 2.56 $M_{\odot}$, with the bimodal mean peaks at 1.351 $M_{\odot}$ and 1.816 $M_{\odot}$ as found and described in details in Rocha et al. (2023) (see Table 3). The mass and spin priors for BHs remain uninformed and the NSs pins are randomly oriented.
@@ 3.3 Models motivated by GW observation of black hole spins
To date, only a small number of NSBH or BHNS binary mergers have been observed Abbott et al. (2023). This makes it difficult to use the observed properties of these binaries to draw strong inferences about their population properties to be used as prior beliefs when interpreting GW230529. However, during the first three LIGO-Virgo-KAGRA (LVK) observing runs, close to a hundred BBH have been observed and the detailed properties of the BH population have been inferred Abbott et al. (2023). The inferred mass of the more massive component of GW230529 lies outside of the observed BH mass distribution and therefore the existing population mass distribution cannot be used. However, we can use the inferred spin distributions inferred from the BBH population as a proxy for the BH spin distribution in NSBH/BHNS binaries. The inferred spin distribution of BH from observations through GWTC-3 is shown in figure 15 of Abbott et al. (2023). The BH spin is modelled through the amplitude $\chi$ and orientation $\cos{\theta}$ of the spin relative to the orbital angular momentum.
@@ 4 Results
Figure 2: The inferred masses and spin projected along the orbital angular momentum $\chi_{z}$ of the components of GW230529, where the subscripts $1$ and $2$, respectively, label the heavier and lighter component of the binary, under five different choices of prior. The LVK prior is an uninformative prior which is flat in redshifted masses and uniform in spin magnitude/orientation; the astrophysical FDT500 prior model is discussed in Section 3.1; the BBH prior uses the $\chi_{1z}$ distribution from the observed BBH population; the pulsar prior uses masses and spin magnitudes from observed pulsars and the pulsar $+$ BBH uses both pulsar and BBH observations. Further details of the models are provided in Section 3 and summarized in Table 1. We note that the $\chi_{1z}$ prior for the astrophysical models is so sharply peaked close to 0 that we have included samples within $<10^{-2}$ to ensure sufficient prior support for re-weighting.
@@
The inferred NS and BH masses based on the astrophysical model are sharply peaked at $1.3^{+0.1}_{-0.1}\,M_{\odot}$ and $4.1^{+0.2}_{-0.3}\,M_{\odot}$ respectively. This renders the secondary clearly a NS with a mass very compatible with the galactic population and the primary a likely BH in the mass gap. While the object is still in the putative “lower mass gap", it is towards the upper edge, the supernova explosion does not have to be extremely rapid (and hence energetic) to achieve such mass. Interestingly, the restriction on the binary masses does not arise due to the astrophysical mass distribution, but rather the spin distribution. As can be seen in Figure 1, the NSBH mass distribution for both NS and BH is rather broad, with support from $1.2-2.5\,M_{\odot}$ for NS and $2.5-6\,M_{\odot}$ for BH. The restriction to (close to) zero spins for both components constrains the mass ratio significantly and, consequently, the component masses. To re-iterate, the BHs of the BHNSs, are always expected to have zero spin, and the NS, being non-recycled, also effectively zero spins. For NSBHs, pulsar recycling extends the NS spin distribution to larger values, although by the time of merger, it still spins down to $\lesssim 0.1-0.2$. The BHs of the NSBHs, have uncertainty associated with their spins — on one hand we model the optimistic efficient tidal spin-up model which gives spins $\chi_{z}\gtrsim 0.5$ (the FDT500_Q model in Fig.1), on the other, we assume them to be non-spinning (FDT500). While the tidal spin-up model can be rejected based upon the observed spins, under non-efficient tides we cannot comment definitively on formation order, i.e. NSBH vs BHNS, in this scenario. While NSBHs certainly prefers a more symmetric mass ratio and lower BH masses, the distribution of both sub-systems is sufficiently broad in mass and peaked near zero spins to be fully consistent with the observation of GW230529.
@@ 5 Discussion
The first, key, point is that the inferred masses and spins of the components of the binary depend critically upon the mass and spin priors used in the analysis and, in particular, the inferred mass distribution is highly dependent upon the spin prior. The fact that the distributions are so reliant on the prior demonstrates that the uncertainties in the observations for this system are large, due to the relatively low SNR of the event. Therefore, we cannot draw strong conclusions about the origin of this event.

### v2

@@ Abstract
We investigate the effects of prior selection on the inferred mass and spin parameters of the neutron star-black hole merger GW230529_181500. Specifically, we explore models motivated by astrophysical considerations, including massive binary and pulsar evolution. We examine mass and spin distributions of neutron stars constrained by radio pulsar observations, alongside black hole spin observations from previous gravitational wave detections. We show that the inferred mass distribution highly depends upon the spin prior. Specifically, under the most restrictive, binary stellar evolution models, we obtain narrower distributions of masses with a black hole mass of $4.3^{+0.1}_{-0.1}\,M_{\odot}$and neutron star mass of $1.3^{+0.03}_{-0.03}\,M_{\odot}$ where, somewhat surprisingly, it is the prior on component spins which has the greatest impact on the inferred mass distributions. Re-weighting using neutron star mass and spin priors from observations of radio pulsars, with black hole spins from observations of gravitational waves, yields the black hole and the neutron star masses to be $3.8^{+0.5}_{-0.6}\,M_{\odot}$ and $1.4^{+0.2}_{-0.1}\,M_{\odot}$ respectively. The sequence of compact object formation — whether the neutron star or the black hole formed first — cannot be determined at the observed signal-to-noise ratio. However, there is no evidence that the black hole was tidally spun up.
@@ 1 Introduction
The gravitational wave (GW) event GW230529_181500 (hereafter abbreviated GW230529) was detected by the LIGO Livingston Observatory during the first part of the fourth observing run of the LIGO-Virgo-KAGRA collaboration (Abac et al., 2024). The binary components are a neutron star and another compact object whose mass likely falls within the “lower mass gap” between the heaviest neutron stars and the lightest black holes, which was previously thought to exist between $2.5\,M_{\odot}$ to approximately $5\,M_{\odot}$ (Bailyn et al., 1998; Özel et al., 2010; Farr et al., 2011). Moreover, the event marks the third confident detection of a black hole-neutron star binary through gravitational waves (Abbott et al., 2021). It followed a probable observation of a pulsar-(mass gap) black hole in the globular cluster NGC 1851 (Barr et al., 2024). Theoretical astrophysics of massive binary evolution has speculated about the existence of black hole-neutron star binaries (O’Shaughnessy et al., 2008; Fryer et al., 2012; Chattopadhyay et al., 2021; Broekgaarden et al., 2021). Still, successive detections of such objects spark more debates on their mass-spin distributions to formation channels (Chandra et al., 2024) to possible electromagnetic counterparts (Zhu et al., 2024; Ronchini et al., 2024).
@@ 2 Incorporating Astrophysical Prior Beliefs
where $\mathbf{\hat{L}}$ is the unit vector in the direction of the orbital angular momentum, and the spin, $\boldsymbol{\chi}_{i}$, is defined as $\boldsymbol{\chi}_{i}=\frac{\mathbf{S}_{i}}{m_{i}^{2}}$, and the normalization is chosen such that a maximally spinning BH has $|\boldsymbol{\chi}|=1$. Consequently, the inferred values of the component masses and spins from a GW observation correlate. Furthermore, the choice of priors can impact the inferred properties for a low signal to noise ratio (SNR) signal such as GW230529. Indeed, due to correlations between the inferred mass ratio and effective spin, choices of spin priors can impact the inferred masses and vice versa.
The gravitational-wave estimation of these parameters for GW230529, as presented in Abac et al. (2024), was performed with minimally informed priors on the mass and spin distributions of the components of the binary, in addition to the location and orientation parameters (Veitch et al., 2015). Specifically, the analysis used mass priors that are flat in the redshifted component masses, within a prescribed range of masses and mass ratios, and uniform in spin magnitude and orientation (see Appendix D of Abac et al. (2024) for details). The output of the original analysis is a set of samples from the posterior probability distribution of the parameters (LIGO Scientific, Virgo, KAGRA collaboration, 2024) [^2] Specifically, we make use of the Combined_PHM_highSpin results for the majority our our analyses. When re-weighting using the astrophysical models described in Section 3.1, which restrict component spins to be close to zero, we use IMRPhenomPv2_NRTidalv2_lowSpin to ensure that we have sufficient samples following re-weighting . We then apply weights to these samples in the ratio of our desired astrophysical before the original, uniform prior. Given an astrophysically motivated prior $\pi_{A}$, the posterior distribution for the parameters $\boldsymbol{\theta}$ describing the progenitor of GW230529 is given by
@@
When performing prior re-weighting, however, it is important that the proposed prior provides support in the same part of the parameter space as the target prior, otherwise the re-weighting will be highly inefficient. Since the original priors are near-uniform in the parameters of interest and also cover a broad parameter range, the astrophysical priors generally comprise a subset of the original ranges. However, in cases where the astrophysical priors are sharply peaked, the re-weighting procedure summarized above can lead to a low number of samples in the astrophysical posterior. In our analyses we ensured a minimum of one thousand effective samples (Kish, 1995; Víctor Elvira, 2018). For the results presented in Section 4, we use those samples to create one-dimensional Kernel Density Estimators of the inferred masses and spins. As 90% credible intervals are the relevant numbers for the scientific conclusions of this work, we checked via bootstrapping tests to what precision the one-dimensional credible intervals could be quoted, and used that precision.
@@ 3.2 Models motivated by radio pulsar observations in star clusters
As an alternative approach to modelling isolated binary evolution, here we utilize the observed population of Galactic globular cluster radio millisecond pulsar [^3] Over a third of all observed millisecond pulsars appear in Galactic globular clusters, which collectively account for less than 0.05% of the total number of stars in the Milky Way. These millisecond pulsars, have spins up to $\chi\simeq$0.4, several times or even order(s) of magnitude lower than normal pulsars. Although NS-BH mergers in star clusters are rare (Ye et al., 2020), and NS s spin down significantly over time before the merger, these high-spin pulsars suggest an alternative formation channel that merits exploration. spin distribution from the Australia Telescope National Facility (ATNF) pulsar catalogue [^4] https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester et al., 2005). From the catalogue, we obtain the pulsar’s spin $P$, and get
@@
where, $c$ is the velocity of light, $G$ is the universal constant of gravitation, $M$ is the mass of the pulsar (assumed to be 1.4 $M_{\odot}$ here) and $I$ is the moment of inertia of the pulsar given by
$I=0.237M_{\mathrm{NS}}R_{\mathrm{NS}}^{2}\bigg{[}1+4.2\frac{M_{\mathrm{NS}}}{M_{\odot}}\frac{\mathrm{km}}{R_{\mathrm{NS}}}+90\Big{(}\frac{M_{\mathrm{NS}}}{M_{\odot}}\frac{\mathrm{km}}{R_{\mathrm{NS}}}\Big{)}^{4}\bigg{]},$ (11)
(computed from Equation 12 of Lattimer & Schutz, 2005, with mass $M_{\mathrm{NS}}$ =1.4 $M_{\odot}$ and radius $R_{\mathrm{NS}}=12$ Km).
Although most pulsars are expected to spin down significantly due to the loss of rotational energy, with decaying magnetic field at the time of merger (Figure 17, 21 of Chattopadhyay et al., 2021), we take the observed NS spin distribution as an upper limit. The maximum observed spin $\chi_{\mathrm{NS}}$ is 0.4, with the primary peak at 0.13 and another at about 0.02. The mass distribution for NSs, while dependent on the NS equation of state and the assumed Tolman-Oppenheimer-Volkoff limit for theoretical studies, can also be estimated from pulsar observations (with the inclusion of radio selection effects). Even with the limited dataset due to the difficulty in measuring NS mass, there have been multiple studies to determine the shape and range of the mass distribution of NSs (Antoniadis et al., 2016; Alsing et al., 2018; Rocha et al., 2023). For this analysis, we assume the NS mass distribution to be a double Gaussian with a maximum mass of 2.56 $M_{\odot}$, with the bimodal mean peaks at 1.351 $M_{\odot}$ and 1.816 $M_{\odot}$ as found and described in details in Rocha et al. (2023) (see Table 3). The mass and spin priors for BHs remain uninformed and the NSs pins are randomly oriented.
@@ 3.3 Models motivated by GW observation of black hole spins
To date, only a small number of NSBH or BHNS binary mergers have been observed (Abbott et al., 2023). This makes it difficult to use the observed properties of these binaries to draw strong inferences about their population properties to be used as prior beliefs when interpreting GW230529. However, during the first three LIGO-Virgo-KAGRA (LVK) observing runs, close to a hundred BBH have been observed and the detailed properties of the BH population have been inferred Abbott et al. (2023). The inferred mass of the more massive component of GW230529 lies outside of the observed BH mass distribution and therefore the existing population mass distribution cannot be used. However, we can use the inferred spin distributions inferred from the BBH population as a proxy for the BH spin distribution in NSBH/BHNS binaries. The inferred spin distribution of BH from observations through GWTC-3 is shown in figure 15 of Abbott et al. (2023). The BH spin is modelled through the amplitude $\chi$ and orientation $\cos{\theta}$ of the spin relative to the orbital angular momentum.
@@ 4 Results
Figure 2: The inferred masses and spin projected along the orbital angular momentum $\chi_{z}$ of the components of GWFch230529, where the subscripts $1$ and $2$, respectively, label the heavier and lighter component of the binary, under five different choices of prior. The LVK prior is an uninformative prior which is flat in redshifted masses and uniform in spin magnitude/orientation; the astrophysical FDT500 prior model is discussed in Section 3.1; the BBH prior uses the $\chi_{1z}$ distribution from the observed BBH population; the pulsar prior uses masses and spin magnitudes from observed pulsars and the pulsar $+$ BBH uses both pulsar and BBH observations. Further details of the models are provided in Section 3 and summarized in Table 1. We note that the $\chi_{1z}$ prior for the astrophysical models is so sharply peaked close to 0 that we have included samples within $<10^{-2}$ to ensure sufficient prior support for re-weighting.
@@
The inferred NS and BH masses based on the astrophysical model are sharply peaked at $1.3^{+0.03}_{-0.03}\,M_{\odot}$ and $4.3^{+0.1}_{-0.1}\,M_{\odot}$ respectively. This renders the secondary clearly a NS with a mass very compatible with the galactic population and the primary a likely BH in the mass gap. While the object is still in the putative “lower mass gap", it is towards the upper edge, the supernova explosion does not have to be extremely rapid (and hence energetic) to achieve such mass. Interestingly, the restriction on the binary masses does not arise due to the astrophysical mass distribution, but rather the spin distribution. As can be seen in Figure 1, the NSBH mass distribution for both NS and BH is rather broad, with support from $1.2-2.5\,M_{\odot}$ for NS and $2.5-6\,M_{\odot}$ for BH. The restriction to (close to) zero spins for both components constrains the mass ratio significantly and, consequently, the component masses. To re-iterate, the BHs of the BHNSs, are always expected to have zero spin, and the NS, being non-recycled, also effectively zero spins. For NSBHs, pulsar recycling extends the NS spin distribution to larger values, although by the time of merger, it still spins down to $\lesssim 0.1-0.2$. The BHs of the NSBHs, have uncertainty associated with their spins — on one hand we model the optimistic efficient tidal spin-up model which gives spins $\chi_{z}\gtrsim 0.5$ (the FDT500_Q model in Fig.1), on the other, we assume them to be non-spinning (FDT500). While the tidal spin-up model can be rejected based upon the observed spins, under non-efficient tides we cannot comment definitively on formation order, i.e. NSBH vs BHNS, in this scenario. While NSBHs certainly prefers a more symmetric mass ratio and lower BH masses, the distribution of both sub-systems is sufficiently broad in mass and peaked near zero spins to be fully consistent with the observation of GW230529.
@@ 5 Discussion
The first key-point is that the inferred masses and spins of the components of the binary depend critically upon the mass and spin priors used in the analysis and, in particular, the inferred mass distribution is highly dependent upon the spin prior. The fact that the distributions are so reliant on the prior demonstrates that the uncertainties in the observations for this system are large, due to the relatively low SNR of the event. Therefore, we cannot draw strong conclusions about the origin of this event.